Betty DeRose, Inc. operates two departments, the handling department and
the packaging department. During April, the handling department reported
the following information:

                                           % complete      % complete
                                units         DM           conversion 
work in process, April 1        18,000        38%             71%
units started during April      80,000
work in process, April 30       44,000        28%             74%

The cost of beginning work in process and the costs added during April
were as follows:

                                 DM         Conversion       Total cost
work in process, April 1      $ 64,996       $152,785         $217,781
costs added during April       190,336        232,407          422,743
total costs                    255,332        385,192          640,524

Calculate the total cost of the 18,000 units in beginning work in process
using the FIFO process costing method.
Answer #1

FIFO method

Given,

Total units in opening work in progress = 18000 units

Units started during the year = 80000 units

Total units in closing work in progress = 44000 units

Under FIFO method, the opening inventory of work in progress is first completed. Thus the entire 18000 units were completed.

Out of 80000 units started during the year, 44000 units are in closing work in progress.

Thus out of 80000 units, 36000 units were started and completed during the period.

Calculation of equivalent number of units and cost per equivalent units for material cost

Units at the opening work in progress 18000 * (100 - 38)% 11,160
Units started and completed 36000 * 100% 36,000
Units at the closing work in progress 44000 * 28% 12,320
Total Equivalent units 59,480

Total Material cost incurred during the year = 190,336

Material Cost per equivalent unit = 190336/59480 = $3.2 per unit

Total material cost incurred on opening work in progress during the year = 11,160 units *3.2 per unit = $35,712

Material Cost incurred in the opening = $64,996

Thus total material cost incurred on 18000 units in beginning work in progress = 35712 + 64996 = $100,708

Calculation of equivalent number of units and cost per equivalent units for conversion cost

Units at the opening work in progress 18000 * (100 - 71)% 5,220
Units started and completed 36000 * 100% 36,000
Units at the closing work in progress 44000 * 74% 32,560
Total Equivalent units 73,780

Total Conversion cost incurred during the year = 232,407

Conversion Cost per equivalent unit = 232407/73780 = $3.15 per unit

Total Conversion cost incurred on opening work in progress during the year = 5,220 units *3.15 per unit = $16,443

Conversion Cost incurred in the opening = $152,785

Thus total Conversion cost incurred on 18000 units in beginning work in progress = 16,443 + 152,785 = $169,228

Therefore total cost incurred on 18000 units in beginning work in progress = $100,708 + $169,228 = $269,936

Kindly Up Vote